(ii) Five years ago, Nuri was thrice as old as Sonu. Ten years later, Nuri will be twice as old as Sonu. How old are Nuri and Sonu?

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Solution: Let us assume, present age of Nuri is x And present age of Sonu is y. According to the given condition, we can write as; x – 5 = 3(y – 5) x – 3y = -10…………………………………..(1) Now, x + 10 = 2(y +10) x – 2y = 10…………………………………….(2) Subtract eq. 1 from 2, to get, y = 20 ………………………………………….(3) Substituting the value of y in eq.1, we get, x – 3.20 = -10 x – 60 = -10 x = 50 Therefore, Age of Nuri is 50 years Age of Sonu is 20 years.